# Dual-Unitary Classical Shadow Tomography

### v1

@@ II.1.3 Set up
    As long as $P_{1}\neq P_{2}$, there exists a local basis rotation $V$ to flip $P_{1}$ as $VP_{1}V^{\dagger}=-P_{1}$ (namely that $P_{1}$ and $V$ anticommute) while keeping $P_{2}$ invariant as $VP_{2}V^{\dagger}=P_{2}$. For example, for qubit Pauli’s, $V=P_{2}$ fulfils this condition since any two different Pauli’s anticommute. Similarly, for $k$-qubit Paulis, pick $V$ to match $P_{2}$ on a site where $P_{1}$ and $P_{2}$ differ, and be identity on the other sites. Under such rotation $V$, the tensor combination $P_{1}\otimes P_{2}$ in Eq. (13) will change sign (unless $P_{1}=P_{2}$ where $V$ cannot be found). Since the integration measure is invariant under any local basis rotations, the integration has to vanish for any $P_{1}\neq P_{2}$ due to its ability to change sign under local basis rotations, hence we must have $P_{1}=P_{2}$ to get a non-vanishing result for the integration. Therefore, integrating over each gate simply produces a sum over more doubled Pauli strings:
@@
    The only Pauli strings that survive the summation in Eq. (17) are $Z$-strings. Therefore, to evaluate the full Pauli weight, we must sum over all possible supports, but weight each non-empty region by $3^{-|b|}$, since that is the fraction of Pauli strings which are $Z$-strings. Therefore, where-ever there is a measurement gate, contract with $(1,1/3)^{T}$. In a standard classical shadow scheme we end with measurements on every site, so that
@@ II.2 Pauli Weight Transfer Matrix
where $b_{i},b_{j}^{\prime}\in\{0,1\}$ and $U$ is integrated over some yet unspecified unitary ensemble that is local-basis invariant. Therefore, the choice of $Z$ is irrelevant and any other Pauli operator would give the same result. With minimal assumptions, we can ascertain the general form of the transfer matrix $\hat{T}_{b,b^{\prime}}$,
@@
The remaining two independent transfer matrix elements $\mathcal{I}_{1},\mathcal{I}_{2}$ parameterize the operator entanglement properties of the unitary ensemble Zanardi (2001) via the Fierz’ identity,
@@
For any two-qubit unitary gate or any two-qubit random unitary ensemble, the parameters $\mathcal{I}_{1},\mathcal{I}_{2}$ can be defined via Eq. (29) (with or without ensemble average). Their feasible domain has been studied in Ref. You and Gu (2018). We simply state the result here: the parameters must be within the box $(\mathcal{I}_{1},\mathcal{I}_{2})\in[0,1]^{\times 2}$ and between the following bounds:
@@
and both bounds are tight (meaning that there exist two-qubit unitary gates along these boundaries). The feasible domain is shown in Fig. 5(a). Interestingly, some well-known two-qubit Clifford gates (and their Pauli-twirled versions) happen to locate at the corners of the region. The random Clifford (or Haar) ensemble corresponds to $\mathcal{I}_{1}=\mathcal{I}_{2}=1/5$, which sits inside the feasible domain. It is worth emphasizing that these bounds can also be understood by replacing the range of $\mathcal{I}_{1},\mathcal{I}_{2}$ with those of the entangling power and gate typicality, respectively, as discussed in Refs. Rather et al. (2020); Brahmachari et al. (2024).
@@ II.3 Dual-unitary case
In the following, we provide an independent explaination of the bound $1/3\leq\mathcal{I}_{2}\leq 1$ for dual-unitary gates without evoking the result cited from Ref. You and Gu (2018). First of all, it’s easy to check that in the dual-unitary case,
@@
which says that the sum of the mutual information that $A$ shares with each other qubit is greater than its entropy. We can prove this equation through a known bound on the tripartite mutual information Yeung (1991). Plugging in the definition for the mutual information on the left-hand side,
@@ II.4 Pauli Weight Dynamics
In this section, we hope to provide some intuition for the Pauli weight dynamics implied by the dual-unitary transfer matrix, Eq. (42). Looking at the total weight one subspace, spanned by $b=10$ and $b=01$, we see that dual-unitarity implies that the diagonal elements vanish, and that the off diagonal elements are always positive and greater than or equal to 1/3. The absence of the diagonal elements implies an interpretation of the dynamics in terms of left and right movers. We may organize the input legs of the transfer matrix (columns of Eq. (42)) so that the left (right) input bit indicates the presence of right (left) mover, and likewise we may organize the output legs of the transfer matrix so that the right (left) output leg corresponds to an outgoing right (left) mover i.e.
@@
Figure 6: The basis states on two adjacent sites. For parity even sites $x+t\cong 0\mod 2$ which support right movers, the basis states ${|0\rangle}$ and ${|1\rangle}$ are represented by the lack of a line and an upward slanting blue line, respectively. For parity odd sites $x+t\cong 1\mod 2$ which support left movers, the basis states ${|0\rangle}$ and ${|1\rangle}$ are represented by the lack of a line and an downward slanting red line, respectively.
We see that dual-unitary dynamics preserve single left/right movers and they can only be absorbed in the presence of the opposite mover. Furthermore, we see that there distinct processes by which the left and right movers can interact: they may branch, fuse, or collide, but there is no process by which a single left-mover and turn into a single-right mover. The rules are summarized in Tab. 1. This is crucially different from the brick-wall Clifford case (with transfer matrix Eq. (44)), where the notion of a left or right mover is more fuzzy, since left movers can turn into right movers at one step. The specific locations of the left-right movers depending on the parity of the lattice position is specific to the staggered brick-wall set up of the circuit (see Fig. 1).
@@
The notion of left/right movers is also useful because it helps us understand the light-cone structure that arises in the case of dual-unitary Pauli weight dynamics. If a left mover is able to avoid being absorbed by any right movers to the left of it, or it is emitted by the left-most right-mover, it will propagate indefinitely until it reaches the system boundaries. The motion of this particle defines the left light-cone boundary. This lightcone structure is intimately related to the fact that dual-unitary brickwork circuits are an example of quantum cellular automata, see for e.g., the excellent review Farrelly (2020).
@@ II.5 Occupation distribution
In this section, we define simple order parameters of the Pauli weight as a distribution to analyze the average occupation over different sub-regions of sites. In addition, we study Monte Carlo simulations to gain some intuition about the Pauli weight dynamics in dual-unitary circuits. We can view the Pauli weight dynamics generated by the transfer matrix in Eq. (42) as a Markov process that defines how a configuration evolves with time. For example, if the configuration on two adjacent sites consists of an incoming left-mover and right-mover, then the outgoing configuration will consist of a single left-mover with probability $\alpha/3$, a single right-mover with probability $\alpha/3$, and both left-mover and right-mover with probability $31-2\alpha/$ 3. Using the rules implied by the transfer matrix, we can generate Pauli weight configurations prior to the final measurement layer. These configurations consist of a $n\times t$ grid, where $n$ is the number of qubits in the Pauli weight Hilbert space and $t$ is the overall circuit depth. Each point in the grid can take value one or zero, depending on if a given site contains a particle or not. These configurations can be generated using Monte Carlo techniques, and some typical configurations are depicted in Fig. 7. Each configuration begins with a single Pauli operator localized at the center of the system $x=n/2$. In these figures, we color the even/odd parity sites differently to more easily differentiate sites which host left-movers from those that host right-movers.
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
There are some interesting features of the dual-unitary case that we would like to bring to the readers’ attention. First, we see that the dual-unitary case (Fig. 7 a-c) inherits a light-cone structure Farrelly (2020). Because we only start with a single-particle at $t=0$, depending on the parity of the initial site, the injection of the this particle will create a forever-propagating excitation that cannot be absorbed by anything (until it hits the boundary). In Fig. 7 a-c, since the particle is always injected on a parity even site, there is always a consistent right light-cone boundary. The first left-mover that branches from this initial particle injection will form a left light-cone boundary also defined by a forever-propagating excitation. When $\alpha$ is small, the probability of emitting this left-mover is small, and so the right-mover mostly propagates freely. When $\alpha$ is large, there is a higher probability of branching and fusing, which creates a condensate of of particles within the light-cone.
@@
In Fig. 8 a, the average occupation $\rho(x,t)$ starting from a single-particle configuration in the dual-unitary case has a clear right boundary due to the initial injection of the right-mover. Because of this, the spatial structure of the average occupation at fixed time-slices are differs markedly from the Clifford case. In particular, because the initial injected right mover cannot be absorbed by a left-mover, the weight on the right light-cone is always one, whereas the weight goes to zero as we approach the left-boundary. This can be seen most acutely in Fig. 9.
Refer to captionRefer to captionRefer to caption
@@
Looking at Fig. 9 a, we see that in the brick-wall Clifford case the average occupation spreads in an unbiased fashion over time, initially localized at near the initial Pauli support. As we evolve in time, the weight becomes more spread out spatially, with the bulk of the weight taking a roughly uniform value of 3/4. This is because deep Clifford circuits eventually thermalize and model global unitaries, which scramble Pauli operators uniformly. Out of all Pauli strings, 3/4 of them have support on site $x$, so the average occupation $\rho(x,t)$ converges is 3/4 inside the effective light-cone.
Contrarily, Fig. 9 b breaks the inversion symmetry in the occupation. Because the initial Pauli weight configuration consists of a single right-mover, this particle will propagate with velocity $v=1$. Furthermore, the location of the right mover always has average occupation one. Therefore, we see a spike at the location of the right light-cone boundary. As we move away from the location of this right-mover, the weight goes to zero. In the dual-unitary case, we see there is much greater fluctuation as we move from even to odd parity sites and vice versa. In Fig. 9 c, we see that that $\alpha$ characterizes how quickly the inside of the light-cone thermalizes. For a fixed time-slice, as we increase $\alpha$, we increase the average value of the weight inside the light-cone. At $t=10$, we see that when $\alpha=2/3$, the weight is uniformly 3/4. However, as we decreases $\alpha$, the weight at the same value of $t$ is smaller.
Refer to captionRefer to captionRefer to caption
@@ II.6 Mean Field Approximation of Pauli Weight Dynamics
According to Eq. (49), the time evolution of $\rho(x,t)$ is determined through the evolution of the many-body Pauli weight state ${|w_{P,t}\rangle}$, which happens through the application of a brick-wall arrangement of transfer matrices Eq. (42) (depicted graphically in Fig. 1b). Since the exact evolution of the many-body state ${|w_{P,t}\rangle}$ is complicated, this does not give us a simple description for the evolution of $\rho(x,t)$. In fact, there is no exact, closed form equation for the evolution for $\rho(x,t)$ because in general the average occupation at a single site will depend on “higher order” contributions from various other average occupation configurations on multiple sites. We instead look for a simplified approach for the evolution of $\rho(x,t)$ inspired by mean-field theory.
One approach is to ignore the correlations between individual sites. Since it is known that the entanglement features has an efficient low-bond dimension description, the same can be expected of the Pauli weight which is entanglement features by a local basis transformation Akhtar et al. (2023). The reason these many-body vectors have low entanglement is because of their positive sign structure Grover and Fisher (2015); Akhtar et al. (2023). We can specify the $k$-th order approximation, $k\leq n$, by saying that all $l$-particle occupations, $l\leq k$, factorize:
@@
We label the variables $m$ and $v$ in such a way to highlight the similarity to the Dirac equation. We see that $m(\rho)$ plays the role of the fermion mass. As the fermion becomes massive, it warps the space around it, which is accomplished through the velocity coefficient $v(\rho)$. If we consider a region of spacetime where $m(\rho)$ is slowly varying, such that we can treat $m(\rho)$ and $v(\rho)$ as constants, the mean field equation matches one of the chiral components of the Dirac field.
@@
with initial boundary condition $\rho(0)=1$. We find that Eq. (57) captures the exact behavior of $\rho$ in the homogeneous case accurately for $t\geq 1$. We find, however, that $\alpha^{\prime}$ becomes a quadratic function of $\alpha$ (see Fig. 12). The quadratic corrections to $\alpha^{\prime}$ arise because the continuum approximation, in which we are replacing the discrete differences with a first derivative, breaks down when $\alpha$ is large. However, the corrections become negligible when $\alpha$ is small, so that $\alpha^{\prime}\rightarrow\alpha$ when $\alpha\rightarrow 0$, as expected.
Refer to captionRefer to caption
@@ II.7 Ramifications for the shadow norm and prediction
where in the second equality, we’ve suppressed the time coordinate because it is assumed that we are considering only the final time-slice before the application of the measurement gates. To simplify our analysis, we invoke the approximation Eq. (50) at $n$-th order: $\rho(A)=\prod_{i\in A}\rho(x)$ for all non-empty regions $A$. This approximation ignores the correlations in the occupation numbers at different sites. While this is a reasonable approximation for the individual occupancies $\rho(A)$, because the Pauli weight vector carries little entanglement due to its limited sign structure Grover and Fisher (2015); Akhtar et al. (2023); Akhtar and You (2020), making the substitution overestimates the Pauli weight. The discrepancy is because the shadow norm involves the sum of an exponential number of occupations, which can contribute in total to a small error in the thermodynamic limit. Performing the substitution gives the right hand side of the following equation:
@@
where $2\leq\beta_{\text{eff}}\leq 3$ and $k\leq k_{\text{eff}}\leq k+2t$ describe the effective base and size of the evolved operator, respectively. The initial decrease in the base for Clifford gates is why shallow depth circuits can provide an advantage over Pauli measurements. For brick-wall Clifford circuits, the overall shadow norm decreases until the optimal depth $t^{*}\sim\log k$, but increasing the circuit depth further actually increases the shadow norm because of $k_{\text{eff}}$ becomes too large Ippoliti et al. (2023).
@@
In general, when $k_{\text{eff}}$ is smaller than the system size $n$, it is better to use brick-wall Clifford circuits for prediction. This is because of the rapid growth of $k_{\text{eff}}$ in dual-unitary circuits, which initially increases the shadow norm more quickly compared to brick-wall Clifford gates, since dual-unitary circuits have maximal butterfly velocity $v_{B}=1$. Therefore, we look instead at operators with full support, thereby eliminating the ballistic dispersion of the occupation as a factor and directly comparing the rates of saturation of dual-unitary and Clifford brick-wall circuits. Consistent with our intuition that dual-unitary circuits are fast-scrambling, we find that brick-wall dual-unitary circuits can out-perform brick-wall Clifford circuits at approximating Clifford measurements and predicting operators with full support.
Refer to captionRefer to caption
Figure 14: We simulate the scaling of the shadow norm on a $n$-qubit Pauli operator, where $n=100$, using MPS simulations of the Pauli weight for a bond dimension of $D=200$, at finite circuit depth $t$. We numerically define $\beta=(\|P\|^{2})^{1/n}$ as how the shadow norm scales with system size. This is represented by the dots in the above figure. The predictions from the mean-field theory Eq. (67) are represented by the dotted lines. (a). We compare scaling of the shadow norm $\beta$ for a fully extensive operator with the predictions from the mean-field theory for various values of $\alpha$. We find strong agreement between the two. (b). We compare the approach to the Clifford measurement limit of dual-unitary circuits to brick-wall Clifford circuits represented by the black line. We find that dual-unitary circuits out can perform brick-wall Clifford circuits for full support operators, approaching the Clifford measurement limit more quickly. For $\alpha=2/3$, there is an advantage at all circuit depths.
@@
The occupation $\rho^{Cl}(t)$ starting from a uniform weight distribution in the Clifford weight was determined by a similar mean field analysis to scale like Ippoliti et al. (2023),

### v2

@@ II.1.3 Set up
    As long as $P_{1}\neq P_{2}$, there exists a local basis rotation $V$ to flip $P_{1}$ as $VP_{1}V^{\dagger}=-P_{1}$ (namely that $P_{1}$ and $V$ anticommute) while keeping $P_{2}$ invariant as $VP_{2}V^{\dagger}=P_{2}$. For example, for qubit Pauli’s, $V=P_{2}$ fulfills this condition since any two different Pauli’s anticommute. Similarly, for $k$-qubit Paulis, pick $V$ to match $P_{2}$ on a site where $P_{1}$ and $P_{2}$ differ, and be the identity operator on all the other sites. Under such rotation $V$, the tensor combination $P_{1}\otimes P_{2}$ in Eq. (13) will change sign (unless $P_{1}=P_{2}$ where $V$ cannot be found). Since the integration measure is invariant under any local basis rotations, the integration has to vanish for any $P_{1}\neq P_{2}$ due to its ability to change sign under local basis rotations, hence we must have $P_{1}=P_{2}$ to get a non-vanishing result for the integration. Therefore, integrating over each gate simply produces a sum over more doubled Pauli strings:
@@
    The only Pauli strings that survive the summation in Eq. (17) are $Z$-strings. Therefore, to evaluate the full Pauli weight, we must sum over all possible supports, but weight each non-empty region by $3^{-|b|}$, since that is the fraction of Pauli strings which are $Z$-strings. Therefore, where-ever there is a measurement gate, contract with $(1,1/3)^{T}$. In a standard classical shadow scheme, we end with measurements on every site, so that
@@ II.2 Pauli Weight Transfer Matrix
where $b_{i},b_{j}^{\prime}\in\{0,1\}$ and $U$ are integrated over some yet unspecified unitary ensemble that is local-basis invariant. Therefore, the choice of $Z$ is irrelevant and any other Pauli operator would give the same result. With minimal assumptions, we can ascertain the general form of the transfer matrix $\hat{T}_{b,b^{\prime}}$,
@@
The remaining two independent transfer matrix elements $\mathcal{I}_{1},\mathcal{I}_{2}$ parameterize the operator entanglement properties of the unitary ensemble [61] via the Fierz’ identity,
@@
For any two-qubit unitary gate or any two-qubit random unitary ensemble, the parameters $\mathcal{I}_{1},\mathcal{I}_{2}$ can be defined via Eq. (29) (with or without ensemble average). Their feasible domain has been studied in Ref. [62]. We simply state the result here: the parameters must be within the box $(\mathcal{I}_{1},\mathcal{I}_{2})\in[0,1]^{\times 2}$ and between the following bounds:
@@
and both bounds are tight (meaning that there exist two-qubit unitary gates along these boundaries). The feasible domain is shown in Fig. 5(a). Interestingly, some well-known two-qubit Clifford gates (and their Pauli-twirled versions) happen to locate at the corners of the region. The random Clifford (or Haar) ensemble corresponds to $\mathcal{I}_{1}=\mathcal{I}_{2}=1/5$, which sits inside the feasible domain. It is worth emphasizing that these bounds can also be understood by replacing the range of $\mathcal{I}_{1},\mathcal{I}_{2}$ with those of the entangling power and gate typicality, respectively, as discussed in Refs. [63, 64].
@@ II.3 Dual-unitary case
In the following, we provide an independent explaination of the bound $1/3\leq\mathcal{I}_{2}\leq 1$ for dual-unitary gates without evoking the result cited from Ref. [62]. First of all, it’s easy to check that in the dual-unitary case,
@@
which says that the sum of the mutual information that $A$ shares with each other qubit is greater than its entropy. We can prove this equation through a known bound on the tripartite mutual information [65]. Plugging in the definition for the mutual information on the left-hand side,
@@ II.4 Pauli Weight Dynamics
In this section, we hope to provide some intuition for the Pauli weight dynamics implied by the dual-unitary transfer matrix, Eq. (42). Looking at the total weight one subspace, spanned by $b=10$ and $b=01$, we see that dual-unitarity implies that the diagonal elements vanish, and that the off-diagonal elements are always positive and greater than or equal to 1/3. The absence of the diagonal elements implies an interpretation of the dynamics in terms of left and right movers. We may organize the input legs of the transfer matrix (columns of Eq. (42)) so that the left (right) input bit indicates the presence of right (left) mover, and likewise we may organize the output legs of the transfer matrix so that the right (left) output leg corresponds to an outgoing right (left) mover i.e.
@@
Figure 6: The basis states on two adjacent sites. For parity even sites $x+t\cong 0\mod 2$ which support right movers, the basis states ${|0\rangle}$ and ${|1\rangle}$ are represented by the lack of a line and an upward slanting blue line, respectively. For parity odd sites $x+t\cong 1\mod 2$ which support left movers, the basis states ${|0\rangle}$ and ${|1\rangle}$ are represented by the lack of a line and a downward slanting red line, respectively.
We see that dual-unitary dynamics preserve single left/right movers and they can only be absorbed in the presence of the opposite mover. Furthermore, we see that there are distinct processes by which the left and right movers can interact: they may branch, fuse, or collide, but there is no process by which a single left-mover and turn into a single-right mover. The rules are summarized in Tab. 1. This is crucially different from the brick-wall Clifford case (with transfer matrix Eq. (44)), where the notion of a left or right mover is more fuzzy, since left movers can turn into right movers at one step. The specific locations of the left-right movers depending on the parity of the lattice position is specific to the staggered brick-wall set up of the circuit (see Fig. 1).
@@
The notion of left/right movers is also useful because it helps us understand the light-cone structure that arises in the case of dual-unitary Pauli weight dynamics. If a left mover is able to avoid being absorbed by any right movers to the left of it, or it is emitted by the left-most right-mover, it will propagate indefinitely until it reaches the system boundaries. The motion of this particle defines the left light-cone boundary. This lightcone structure is intimately related to the fact that dual-unitary brick-wall circuits are an example of quantum cellular automata, see for e.g., the excellent review [66].
@@ II.5 Occupation distribution
In this section, we define simple order parameters of the Pauli weight as a distribution to analyze the average occupation over different sub-regions of sites. In addition, we study Monte Carlo simulations to gain some intuition about the Pauli weight dynamics in dual-unitary circuits. We can view the Pauli weight dynamics generated by the transfer matrix in Eq. (42) as a Markov process that defines how a configuration evolves with time. For example, if the configuration on two adjacent sites consists of an incoming left-mover and right-mover, then the outgoing configuration will consist of a single left-mover with probability $\alpha/3$, a single right-mover with probability $\alpha/3$, and both left-mover and right-mover with probability $31-2\alpha/$ 3. Using the rules implied by the transfer matrix, we can generate Pauli weight configurations prior to the final measurement layer. These configurations consist of a $n\times t$ grid, where $n$ is the number of qubits in the Pauli weight Hilbert space and $t$ is the overall circuit depth. Each point in the grid can take a value of one or zero, depending on if a given site contains a particle or not. These configurations can be generated using Monte Carlo techniques, and some typical configurations are depicted in Fig. 7. Each configuration begins with a single Pauli operator localized at the center of the system $x=n/2$. In these figures, we color the even/odd parity sites differently to more easily differentiate sites which host left-movers from those that host right-movers.
@@
There are some interesting features of the dual-unitary case that we would like to bring to the readers’ attention. First, we see that the dual-unitary case (Fig. 7 a-c) inherits a light-cone structure [66]. Because we only start with a single-particle at $t=0$, depending on the parity of the initial site, the injection of the this particle will create a forever-propagating excitation that cannot be absorbed by anything (until it hits the boundary). In Fig. 7 a-c, since the particle is always injected on a parity even site, there is always a consistent right light-cone boundary. The first left-mover that branches from this initial particle injection will form a left light-cone boundary also defined by a forever-propagating excitation. When $\alpha$ is small, the probability of emitting this left-mover is small, and so the right-mover mostly propagates freely. When $\alpha$ is large, there is a higher probability of branching and fusing, which creates a condensate of of particles within the light-cone.
@@
In Fig. 8 a, the average occupation $\rho(x,t)$ starting from a single-particle configuration in the dual-unitary case has a clear right boundary due to the initial injection of the right-mover. Because of this, the spatial structure of the average occupation at fixed time-slices differs markedly from the Clifford case. In particular, because the initial injected right mover cannot be absorbed by a left-mover, the weight on the right light-cone is always one, whereas the weight goes to zero as we approach the left-boundary. This can be seen most acutely in Fig. 9.
@@
Looking at Fig. 9 a, we see that in the brick-wall Clifford case, the average occupation spreads in an unbiased fashion over time, initially localized at near the initial Pauli support. As we evolve in time, the weight becomes more spread out spatially, with the bulk of the weight taking a roughly uniform value of 3/4. This is because deep Clifford circuits eventually thermalize and model global unitaries, which scramble Pauli operators uniformly. Out of all Pauli strings, 3/4 of them have support on site $x$, so the average occupation $\rho(x,t)$ converges is 3/4 inside the effective light-cone.
Contrarily, Fig. 9 b breaks the inversion symmetry in the occupation. Because the initial Pauli weight configuration consists of a single right-mover, this particle will propagate with velocity $v=1$. Furthermore, the location of the right mover always has an average occupation of one. Therefore, we see a spike at the location of the right light-cone boundary. As we move away from the location of this right-mover, the weight goes to zero. In the dual-unitary case, we see there is much greater fluctuation as we move from even to odd parity sites and vice versa. In Fig. 9 c, we see that that $\alpha$ characterizes how quickly the inside of the light-cone thermalizes. For a fixed time-slice, as we increase $\alpha$, we increase the average value of the weight inside the light-cone. At $t=10$, we see that when $\alpha=2/3$, the weight is uniformly 3/4. However, as we decrease $\alpha$, the weight at the same value of $t$ is smaller.
@@ II.6 Mean Field Approximation of Pauli Weight Dynamics
According to Eq. (49), the time evolution of $\rho(x,t)$ is determined through the evolution of the many-body Pauli weight state ${|w_{P,t}\rangle}$, which happens through the application of a brick-wall arrangement of transfer matrices Eq. (42) (depicted graphically in Fig. 1b). Since the exact evolution of the many-body state ${|w_{P,t}\rangle}$ is complicated, this does not give us a simple description of the evolution of $\rho(x,t)$. In fact, there is no exact, closed-form equation for the evolution for $\rho(x,t)$ because in general the average occupation at a single site will depend on “higher-order” contributions from various other average occupation configurations on multiple sites. We instead look for a simplified approach for the evolution of $\rho(x,t)$ inspired by mean-field theory.
One approach is to ignore the correlations between individual sites. Since it is known that the entanglement features has an efficient low-bond dimension description, the same can be expected of the Pauli weight which is entanglement features by a local basis transformation [53]. The reason these many-body vectors have low entanglement is because of their positive sign structure [67, 48]. We can specify the $k$-th order approximation, $k\leq n$, by saying that all $l$-particle occupations, $l\leq k$, factorize:
@@
We label the variables $m$ and $v$ in such a way to highlight the similarity to the Dirac equation. We see that $m(\rho)$ plays the role of the fermion mass. As the fermion becomes massive, it warps the space around it, which is accomplished through the velocity coefficient $v(\rho)$. If we consider a region of spacetime where $m(\rho)$ is slowly varying, such that we can treat $m(\rho)$ and $v(\rho)$ as constants, the mean-field equation matches one of the chiral components of the Dirac field.
@@
with initial boundary condition $\rho(0)=1$. We find that Eq. (57) captures the exact behavior of $\rho$ in the homogeneous case accurately for $t\geq 1$. We find, however, that $\alpha^{\prime}$ becomes a quadratic function of $\alpha$ (see Fig. 12). The quadratic corrections to $\alpha^{\prime}$ arise because the continuum approximation, in which we are replacing the discrete differences with a first derivative, breaks down when $\alpha$ is large. However, the corrections become negligible when $\alpha$ is small so that $\alpha^{\prime}\rightarrow\alpha$ when $\alpha\rightarrow 0$, as expected.
@@ II.7 Ramifications for the shadow norm and prediction
where in the second equality, we’ve suppressed the time coordinate because it is assumed that we are considering only the final time slice before the application of the measurement gates. To simplify our analysis, we invoke the approximation Eq. (50) at $n$-th order: $\rho(A)=\prod_{i\in A}\rho(x)$ for all non-empty regions $A$. This approximation ignores the correlations in the occupation numbers at different sites. While this is a reasonable approximation for the individual occupancies $\rho(A)$, because the Pauli weight vector carries little entanglement due to its limited sign structure [67, 48, 31], making the substitution overestimates the Pauli weight. The discrepancy is because the shadow norm involves the sum of an exponential number of occupations, which can contribute in total to a small error in the thermodynamic limit. Performing the substitution gives the right hand side of the following equation:
@@
where $2\leq\beta_{\text{eff}}\leq 3$ and $k\leq k_{\text{eff}}\leq k+2t$ describe the effective base and size of the evolved operator, respectively. The initial decrease in the base for Clifford gates is why shallow depth circuits can provide an advantage over Pauli measurements. For brick-wall Clifford circuits, the overall shadow norm decreases until the optimal depth $t^{*}\sim\log k$, but increasing the circuit depth further actually increases the shadow norm because of $k_{\text{eff}}$ becomes too large [51].
@@
Figure 14: We simulate the scaling of the shadow norm on a $n$-qubit Pauli operator, where $n=100$ in (a-b), using MPS simulations of the Pauli weight for a bond dimension of $D=200$, at finite circuit depth $t$. We numerically define $\beta=(\|P\|^{2})^{1/n}$ as how the shadow norm scales with system size. This is represented by the dots in the above figure. The predictions from the mean-field theory Eq. (67) are represented by the dotted lines. (a). We compare the scaling of the shadow norm $\beta$ for a fully extensive operator with the predictions from the mean-field theory for various values of $\alpha$. We find strong agreement between the two. (b). We compare the approach to the Clifford measurement limit of dual-unitary circuits to brick-wall Clifford circuits represented by the black line. We find that dual-unitary circuits out can perform brick-wall Clifford circuits for full support operators, approaching the Clifford measurement limit more quickly. For $\alpha=2/3$, there is an advantage at all circuit depths. (c). We compare brick-wall dual-unitary circuits with brick-wall Clifford gates at finite size for operators with extensive support. The mean-field approximation, which lower bounds the shadow norm in the thermodynamic limit, is represented by the boundary of the shaded region. We find that for small systems, the advantage of using dual-unitary gates over Clifford gates increases. This is because the boundary effects in the dual-unitary case are favorable for dissipating excitations in the occupation, thereby lowering the shadow norm. We also find that the exact numerics converge to the mean-field approximation result above $n\approx 30$, whereas below $n\approx 30$ the mean-field approximation sets a reliable upper bound to the shadow norm.
In general, when $k_{\text{eff}}$ is smaller than the system size $n$, it is better to use brick-wall Clifford circuits for prediction. This is because of the rapid growth of $k_{\text{eff}}$ in dual-unitary circuits, which initially increases the shadow norm more quickly compared to brick-wall Clifford gates, since dual-unitary circuits have maximal butterfly velocity $v_{B}=1$. Therefore, we look instead at operators with full support, thereby eliminating the ballistic dispersion of the occupation as a factor and directly comparing the rates of saturation of dual-unitary and Clifford brick-wall circuits. Consistent with our intuition that dual-unitary circuits are fast-scrambling, we find that brick-wall dual-unitary circuits can outperform brick-wall Clifford circuits at approximating Clifford measurements and predicting operators with full support.
@@
The occupation $\rho^{Cl}(t)$ starting from a uniform weight distribution in the Clifford weight was determined by a similar mean field analysis to scale like [51],